A number of two-digit has 3 for its unit digit and the sum of the digit is 2/11 of the number itself. The number is
Answer by htmentor(1343)

The number can be written as 10t + u, where t is the tens place, u is the units place
Given: u = 3 and sum of digits = 2/11 of the number
(2/11)(10t + 3) = t + 3
Solve for t:
(9/11)t = 33/11 - 6/11 = 27/11
t = 3
Check: (2/11)33 = 2*3 = 6